OpenStudy (anonymous):

Find S8 for the geometric series 3 + -6 + 12 + -24 +…

OpenStudy (anonymous):

-96 -255 768 192

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

What is the common ratio here choco?

OpenStudy (anonymous):

the ratiois 3 again

mathslover (mathslover):

No. See I have , 3 + (-6) + (12) + (-24) ... 3 * -2 = -6 -6 * (-2) = 12 12 * (-2) = -24 That is each number is getting multiplied by (-2) to get the another number of the series. So common ratio here is (-2). Ok?

OpenStudy (anonymous):

OK got it!

mathslover (mathslover):

Do you know the formula for getting S_n of a geometric series (infinite) ?

mathslover (mathslover):

^Not infinite. Its finite to 8.

OpenStudy (anonymous):

no thats the problem that Im having

OpenStudy (anonymous):

you there @mathslover

mathslover (mathslover):

Yes, please wait.

OpenStudy (anonymous):

ok

mathslover (mathslover):

Formula : \(\cfrac{a(1-r^n)}{(1-r)} \)

mathslover (mathslover):

Put r = (-2) , a = 3 and n = 8 . Solve it now.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got 257

mathslover (mathslover):

Try it again. It must be negative .

OpenStudy (anonymous):

ok -255

OpenStudy (anonymous):

?

mathslover (mathslover):

Right!

mathslover (mathslover):

Good work.

OpenStudy (anonymous):

Thank You again!

mathslover (mathslover):

You're welcome and Best of Luck again ;)